Two 1.50-V batteries

Two 1.50-V batteries
Answer #1

a) If R is the resistance of the bulb we can use R = V/I

total resistance of the circuit=0.21+0.16+R=0.37+R

total voltage=2*1.5=3v

current=3/0.37+R

0.37+R=3/600*10^-3=5

R=5-0.37=4.63 ohm
b) total power = V*I = 0.6*3 = 1.8W

Power in the internal resistance = I^2*R = 0.6^2 * 0.37 = 0.1332W

Therefore fraction of total power dissipated in the batteries = 0.1332/ 1.8 = 0.074